OpenStudy (anonymous):

a round cookie divide into 8 equal piece,has adiameter of 16 inches.maty eats 2 pieces.what the combine area of the two pieces that she eats

OpenStudy (anonymous):

area of cookie = pi * r^2 area that he eats = 2/8 of this

OpenStudy (anonymous):

2 pieces eqauls to area of sector with radius 8 and angle of 90 degree because 8 pieces =360 degree hence 2pieces =90 degree Area of sector is ((angle*pi)/360)*8*8